OpenStudy (anonymous):

P(-2,-1) m=2 write an equation of a line

OpenStudy (zzr0ck3r):

so our form is y = mx+b we know that m =2 so we have y = 2x+b now we want to find what b is so we sub in our point for x and y let x = -2 and let y = -1 y=2x+b becomes -1=2(-2)+b can you solve for b?

OpenStudy (anonymous):

b would = 3 -1=2(-2)+b -1=-4+b 3=b

OpenStudy (zzr0ck3r):

yep

OpenStudy (zzr0ck3r):

y = 2x+3

OpenStudy (anonymous):

okay thank you